OpenStudy (anonymous):

Suppose two people begin walking in the same direction at different speeds. Walker A starts at the 0.5 meter mark and walks at a speed of 1 m/s. Walker B starts at the 2-meter mark and walks at a speed of 0.5 m/s. When and where will walker A pass walker B?

OpenStudy (anonymous):

\[x + 0.5 = 0.5x + 2\] \[0.5x = 1.5\] \[x = 3\] \[y = f(3) = (3) + 0.5 = 3.5 \] Both walkers will meet at 3.5 meters after 3 seconds.

OpenStudy (anonymous):

Thank you so much, i had a very hard time with this one, you helped a lot.